OpenStudy (anonymous):

how do i solve this problem that i got wrong??? You write each letter of the word mathematician on a piece of paper and put each letter in a bag.You randomly select one piece from the bag. What is P(t or A)? A.5/13 B.5/8 C.1/3 D.8/13

OpenStudy (anonymous):

@Destinymasha @Shay17

OpenStudy (acxbox22):

M A T H E M A T I C I A N 13 letters probability the letter is a A or T A=3 T=2 3+2=5 5/13 is the probability

OpenStudy (anonymous):

@acxbox22 how would you solve this one becuas ei also got it wrong You have 3 pairs of white socks,4 pairs of black socks,and 2 pairs of colored socks in your drawer. What is the probability that you reach into the drawer and randomly select a pair of white socks? A.1/2 B.1/3 C.2/3 D.3/7

OpenStudy (acxbox22):

B 3/9=1/3